Solve for y in terms of x: absolute value of 3y-2/3x less than 1/2.
Answer by stanbon(75887)

absolute value of 3y-2/3x less than 1/2.
|3y-(2/3)x|<=(1/2) means
-(1/2)<=[3y-(2/3)x]<=1/2
-(1/2)+(2/3)x<=3y<=(1/2)+(2/3)x
-(1/6)+(2/9)x<=y<=(1/6)+(2/9)x
Cheers,
Stan H.